RC meeting and Run Plan, Friday November 11th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time:at 3:30pm today

  • Location: Counting house conference room

Useful Links

Current Status

Accelerator and Hall Status

  • Beam maintenance will take place on Tuesday. MCC will take beam away around 6:45am and the whole procedure will take at least one and half shift.
  • After going though several tuning procedure to recover from frequent beam trip MCC decided to send 43uA to Hall A and 48uA to Hall C. This seems to be the most stable condition for running.
  • Yesterday early swing shift, Hunter did an EPR to get a calibration constant. The result was 1.7
  • We finished Ion chamber calibration on reference cell (H2) around 10 pm yesterday. It didn't seem prudent to start the process of establishing beam on He3 target at the beginning of owl shift, so we kept taking data on H2 overnight.
  • Simona started the process of establishing beam on He3 cell right at the beginning of day shift. It went smoothly and we are taking production data since 9 am.
  • Target polarization kept rising consistantly overnight. Current polarization is 37.8% according to the last NMR sweep (12:36 pm). This is the highest polarization we have seen since the start of GEn!
  • We think Windmill can do even better! So target group is trying to figure out ways to optimize polarization.

Current Activity and Issues

Beamline

  • Dump right Ion chamber unresponsive. For now director of operations agreed to allow running with just one dump IC. - J. Kowal will make an access during maintenance day to fix it (He will also look at the Compton IC).

Target

  • Todd made escorted access to adjust the variacs that control the target oven heater. - Todd
  • One of the laser fiber temperature is higher than usual. Is it safe to run like that? - Target group decided to discuss offline. Any update?
  • Calibration constant at the current kinematic setting. - Hunter will do EPR measurement to get the calibration constant as soon as beam checkout is done. Result: 1.7

Subsystems Updates

Polarized Target

  • Current Polarization: 37.8% (12:36pm)

Any additional subsystem updates?

Run Plan

Current Plan

  • Continue taking production data
  1. 43uA of beam current | Use ps3=ps5=0, ps1=9, ps2=15, rest =-1 | One hour long runs
  2. Move the target to the pickup coil position and do an NMR sweep every 2/3 hours
  3. Return to step 1.

Requests

  • Moller measurement - Schedule it on Thursday day shift
  • Optics runs with sieve in
  • GEM luminosity study

Instructions for Shift Crew

  • Keep an eye out for drifting BPM positions
  • If beam has been down due to accelerator issues perform a checkout of the beam profile by doing harp scans and check beam centering on target using the Carbon hole target, see Simona's procedure: procedure, If in doubt about whether you should go through the beam checkout procedure, call this number: 7578693131.
  • Cross check the rate with what is expected: He3, H2
  • Cross check with the expected ion chamber readings: log
  • If holding field is changed for any reason, do NOT perform an NMR measurement.

Be sure to follow these procedures:

Miscellaneous

Experimental progress (for MCC planning meeting)

  • 6.1% of accumulated charge for production data (13.55 Coulombs out of 232)
  • 32% Experimental run time (47 out of 141 days)